# Existence of non-Abelian vortices in a coupled 4D–2D quantum field theory

### v1

@@ 4.1 Existence and uniqueness of the critical point
Next we need to estimate the term $M_{i}(u_{i})(i=1,2)$ on the right-hand side of the above. To proceed further, as in [16], we can choose a decomposition $u_{i}=u_{i+}-u_{i-}(i=1,2)$ with $u_{+}=\max\{0,u\}$ and $u_{-}=\max\{0,-u\}$ for $u\in\mathbb{R}$. Then $M_{i}(u_{i})=M_{i}(u_{i+})+M_{i}(-u_{i-})(i=1,2)$.
According to the elementary inequality $\mathrm{e}^{t}-1\geq t$ for $t\in\mathbb{R}$, we have $\mathrm{e}^{2u_{1+}}-1\geq 2u_{1+}$, which yield the lower bound
$M_{1}(u_{1+})=\frac{2}{N-1}\int_{\mathbb{R}^{2}}u_{1+}\left(\mathrm{e}^{2u_{1+}}-1\right)\mathrm{d}x$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\geq\frac{1}{N-1}\left\|2u_{1+}\right\|_{2}^{2}>\frac{1}{N-1}\int_{\mathbb{R}^{2}}\frac{\left(2u_{1+}\right)^{2}}{1+2u_{1+}}\mathrm{d}x.$ (4.20)
Then applying the inequality $1-\mathrm{e}^{-t}\geq\frac{t}{1+t}$ for $t\geq 0$, we can estimate $M_{1}(-u_{1-})$ as follows
$M_{1}(-u_{1-})=\frac{2}{N-1}\int_{\mathbb{R}^{2}}u_{1-}\left(1-\mathrm{e}^{-2u_{1-}}\right)\mathrm{d}x$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\geq\frac{1}{N-1}\int_{\mathbb{R}^{2}}\frac{\left(2u_{1-}\right)^{2}}{1+2u_{1-}}\mathrm{d}x.$ (4.21)
@@
Recall the lower estimate for $M_{1}(u_{1+})$ obtained earlier. As a consequence, we find that
$M_{1}(u_{1})\geq\frac{1}{N-1}\int_{\mathbb{R}^{2}}\frac{\left(2u_{1}\right)^{2}}{1+|2u_{1}|}\mathrm{d}x.$ (4.22)
Analogous estimates could be made for $M_{2}(u_{2})$
$M_{2}(u_{2})\geq\frac{1}{N-1}\int_{\mathbb{R}^{2}}\frac{\left(2u_{2}\right)^{2}}{1+|2u_{2}|}\mathrm{d}x.$ (4.23)
@@
$DI(w_{1},w_{2})(w_{1},w_{2})-\frac{20N}{9(3N-1)(N-1)}\int_{\mathbb{R}^{2}}\left(\left|\nabla u_{1}\right|^{2}+\left|\nabla u_{2}\right|^{2}\right)\mathrm{d}x$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\geq\frac{1}{N-1}\int_{\mathbb{R}^{2}}\left\{\frac{\left(2u_{1}\right)^{2}}{1+|2u_{1}|}+\frac{\left(2u_{2}\right)^{2}}{1+|2u_{2}|}\right\}\mathrm{d}x.$ (4.24)
In order to analyze $\left\|2u_{i}\right\|_{L^{2}(\mathbb{R}^{2})}(i=1,2)$, we now stress the Gagliardo–Nirenberg–Sobolev embedding inequality[12] over $W^{1,2}(\mathbb{R}^{2})$
@@
$\left(\int_{\mathbb{R}^{2}}(2u_{i})^{2}\mathrm{d}x\right)^{2}=\left(\int_{\mathbb{R}^{2}}\frac{|2u_{i}|}{1+|2u_{i}|}\left(1+|2u_{i}|\right)|2u_{i}|\mathrm{d}x\right)^{2}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\leq\int_{\mathbb{R}^{2}}\frac{\left(2u_{i}\right)^{2}}{\left(1+|2u_{i}|\right)^{2}}\mathrm{d}x\int_{\mathbb{R}^{2}}\left(1+|2u_{i}|\right)^{2}|2u_{i}|^{2}\mathrm{d}x$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\leq 4\int_{\mathbb{R}^{2}}\frac{\left(2u_{i}\right)^{2}}{\left(1+|2u_{i}|\right)^{2}}\mathrm{d}x\int_{\mathbb{R}^{2}}|2u_{i}|^{2}\mathrm{d}x\left(1+\int_{\mathbb{R}^{2}}|2\nabla u_{i}|^{2}\mathrm{d}x\right)$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\leq\frac{1}{2}\left(\int_{\mathbb{R}^{2}}|2u_{i}|^{2}\mathrm{d}x\right)^{2}+32\Bigg{\{}1+\left[\int_{\mathbb{R}^{2}}\frac{\left(2u_{i}\right)^{2}}{\left(1+|2u_{i}|\right)^{2}}\mathrm{d}x\right]^{4}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\left(\int_{\mathbb{R}^{2}}|2\nabla u_{i}|^{2}\mathrm{d}x\right)^{4}\Bigg{\}},~{}~{}i=1,2$ (4.26)
@@
$\left\|2u_{i}\right\|_{L^{2}(\mathbb{R}^{2})}\leq C\left[1+\int_{\mathbb{R}^{2}}\frac{\left(2u_{i}\right)^{2}}{\left(1+|2u_{i}|\right)^{2}}\mathrm{d}x+\int_{\mathbb{R}^{2}}|2\nabla u_{i}|^{2}\mathrm{d}x\right],~{}~{}i=1,2,$ (4.27)
@@
$F\left(w_{1},w_{2}\right)=\frac{1}{N-1}\left(\mathrm{e}^{2w_{1}}-1\right)+\mathrm{e}^{2\left(\frac{2N-1}{3N-1}w_{1}+w_{2}\right)}-1-\frac{4N^{2}}{(3N-1)(N-1)}w_{1}-2w_{2},\$
@@ 4.2 Asymptotic behavior and quantized integrals
Finally, we estimate the decay rates for $u_{1},u_{2}$ and $|\nabla(mu_{1}+nu_{2})|,|\nabla(pu_{1}+qu_{2})|$, where $m,n,p,q$ are as defined by (3.20). For the coefficient matrix $A$ of the system (4.2), there exists a diagonal matrix $B$ such that the matrix $BA=M$ is positive definite and symmetric, where
@@
Therefore, we reduce (3.11)–(3.12) into
@@
$\triangle g^{2}\geq\lambda_{0}g^{2}-C(x)g^{2},~{}~{}x\in\mathbb{R}^{2},$ (4.43)
@@
$\triangle g^{2}\geq\left(1-\frac{\varepsilon}{2}\right)\lambda_{0}g^{2},~{}~{}x\in\mathbb{R}^{2}.$ (4.44)
@@
$~{}~{}~{}~{}\,\triangleq 2P\mathbf{h}+2P\left(E(x)-I\right)\mathbf{h}.$ (4.47)
@@
such that $TPT^{-1}=\Lambda=\mathrm{diag}\{\lambda_{3},\lambda_{4}\}$, where $\alpha=\dfrac{3}{2}-\dfrac{1}{2N}$ and
@@
$\triangle\mathbf{H}=2\Lambda\mathbf{H}+2TP\left(E(x)-I\right)T^{-1}\mathbf{H}.$ (4.49)
@@
$~{}~{}~{}~{}~{}~{}=4\mathbf{H}^{\tau}\Lambda\mathbf{H}+4\mathbf{H}^{\tau}TP\left(E(x)-I\right)T^{-1}\mathbf{H}$
$~{}~{}~{}~{}~{}~{}\geq\lambda\mathbf{|H|}^{2}-b(x)\mathbf{|H|}^{2},~{}~{}x\in\mathbb{R}^{2},$ (4.50)
@@
We next are in a position to calculate the quantized integrals applying the decay estimates. Indeed, we have obtained the unique solution $(u_{1},u_{2})$ of the system (3.15) subject to the boundary condition (3.13) which vanish at infinity exponentially fast. Additionally, we deduce that $|\nabla(mu_{1}+nu_{2})|,|\nabla(pu_{1}+qu_{2})|$ all vanish at infinity at least as fast as $|x|^{-3}$, where $m,n,p,q$ are as defined by (3.20). Thus, according to the divergence theorem, we see that
$\int_{\mathbb{R}^{2}}\triangle\left(mu_{1}(x)+nu_{2}(x)\right)\mathrm{d}x=\int_{\mathbb{R}^{2}}\triangle\left(pu_{1}(x)+qu_{2}(x)\right)\mathrm{d}x=0.$ (4.53)
Now according to (4.53), (3.11) and (3.12), we can get the quantized integrals (3.1) as stated in Theorem 3.1.

### v2

@@ 4.1 Existence and uniqueness of the critical point
Next we need to estimate the term $M_{i}(P_{i})(i=1,2)$ on the right-hand side of the above. To proceed further, as in [16], we can choose a decomposition $P_{i}=P_{i+}-P_{i-}(i=1,2)$ with $P_{+}=\max\{0,P\}$ and $P_{-}=\max\{0,-P\}$ for $P\in\mathbb{R}$. Then $M_{i}(P_{i})=M_{i}(P_{i+})+M_{i}(-P_{i-})$.
According to the elementary inequality $\mathrm{e}^{t}-1\geq t$ for $t\in\mathbb{R}$, we can obtain
$M_{1}(P_{1+})\geq\int_{\mathbb{R}^{2}}(2P_{1+})^{2}\mathrm{d}x+\int_{\mathbb{R}^{2}}2P_{1+}\left(2u_{1}^{0}+X_{1}\right)\mathrm{d}x$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\geq\frac{1}{2}\left\|2P_{1+}\right\|_{2}^{2}-\frac{1}{2}\int_{\mathbb{R}^{2}}\left(2u_{1}^{0}+X_{1}\right)^{2}\mathrm{d}x$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\geq\frac{1}{2}\int_{\mathbb{R}^{2}}\frac{\left(2P_{1+}\right)^{2}}{1+2P_{1+}}\mathrm{d}x-C,$ (4.24)
where we have used the fact $u_{1}^{0},X_{1}\in L^{2}(\mathbb{R}^{2})$. Here and what follows we use $C$ to denote a general positive constant that can take different values at different places. Then applying the inequality
$1-\mathrm{e}^{-t}\geq\frac{t}{1+t},~{}~{}~{}\forall t\geq 0,\$
we can estimate $M_{1}(-P_{1-})$ as follows
$M_{1}(-P_{1-})=\int_{\mathbb{R}^{2}}2P_{1-}\left(1-\mathrm{e}^{2(u_{1}^{0}-P_{1-})}-X_{1}\right)\mathrm{d}x$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\,~{}=\int_{\mathbb{R}^{2}}2P_{1-}\left[\mathrm{e}^{2u_{1}^{0}}\left(1-\mathrm{e}^{-2P_{1-}}\right)-\mathrm{e}^{2u_{1}^{0}}+1-X_{1}\right]\mathrm{d}x$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\,~{}\geq\int_{\mathbb{R}^{2}}2P_{1-}\left(\mathrm{e}^{2u_{1}^{0}}\frac{2P_{1-}}{1+2P_{1-}}-\mathrm{e}^{2u_{1}^{0}}+1-X_{1}\right)\mathrm{d}x$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\,~{}=\int_{\mathbb{R}^{2}}\frac{2P_{1-}}{1+2P_{1-}}\left[\left(1+2P_{1-}\right)\left(1-X_{1}-\mathrm{e}^{2u_{1}^{0}}\right)+2P_{1-}\mathrm{e}^{2u_{1}^{0}}\right]\mathrm{d}x$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\,~{}=\int_{\mathbb{R}^{2}}\frac{2P_{1-}}{1+2P_{1-}}\left(1-X_{1}-\mathrm{e}^{2u_{1}^{0}}\right)\mathrm{d}x+\int_{\mathbb{R}^{2}}\frac{\left(2P_{1-}\right)^{2}}{1+2P_{1-}}\left(1-X_{1}\right)\mathrm{d}x.$ (4.25)
Noting the definition of (4.2), then we can choose $\tau>0$ sufficiently large such that
$X_{1}(x)<\frac{1}{2},~{}~{}~{}\forall x\in\mathbb{R}^{2}.\$
In view of the fact $1-\mathrm{e}^{2u_{1}^{0}}$ and $X_{1}$ both belong to $L^{2}(\mathbb{R}^{2})$, we see that
$\int_{\mathbb{R}^{2}}\frac{2P_{1-}}{1+2P_{1-}}\left|1-X_{1}-\mathrm{e}^{2u_{1}^{0}}\right|\mathrm{d}x\leq\varepsilon\int_{\mathbb{R}^{2}}\frac{\left(2P_{1-}\right)^{2}}{(1+2P_{1-})^{2}}\mathrm{d}x+C(\varepsilon),$ (4.26)
where $\varepsilon>0$ could be taken to be arbitrarily small. Therefore, combining (4.26) and (4.1), we find that
$M_{1}(-P_{1-})\geq\frac{1}{4}\int_{\mathbb{R}^{2}}\frac{\left(2P_{1-}\right)^{2}}{(1+2P_{1-})^{2}}\mathrm{d}x-C_{1}(\varepsilon),$ (4.27)
@@
provided that $\varepsilon<\frac{1}{4}$. Recall the lower estimate for $M_{1}(P_{1+})$ obtained earlier. As a consequence, we find that
$M_{1}(P_{1})\geq\frac{1}{4}\int_{\mathbb{R}^{2}}\frac{\left(2P_{1}\right)^{2}}{(1+|2P_{1}|)^{2}}\mathrm{d}x-C.$ (4.28)
Analogous estimates could be made for $M_{2}(P_{2})$
$M_{2}(P_{2})\geq\frac{1}{4}\int_{\mathbb{R}^{2}}\frac{\left(2P_{2}\right)^{2}}{(1+|2P_{2}|)^{2}}\mathrm{d}x-C.$ (4.29)
@@
$DI(w_{1},w_{2})(w_{1},w_{2})-\frac{5}{9}\min\left\{\frac{2\alpha-1}{2\beta},\frac{4\alpha}{\alpha+\beta}\right\}\int_{\mathbb{R}^{2}}\left(\left|\nabla P_{1}\right|^{2}+\left|\nabla P_{2}\right|^{2}\right)\mathrm{d}x$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\geq\frac{2\alpha-1}{8\beta}\int_{\mathbb{R}^{2}}\frac{\left(2P_{1}\right)^{2}}{(1+|2P_{1}|)^{2}}\mathrm{d}x+\frac{1}{4}\int_{\mathbb{R}^{2}}\frac{\left(2P_{2}\right)^{2}}{(1+|2P_{2}|)^{2}}\mathrm{d}x-C.$ (4.30)
In order to analyze $\left\|2P_{i}\right\|_{L^{2}(\mathbb{R}^{2})}(i=1,2)$, we now stress the Gagliardo–Nirenberg–Sobolev embedding inequality[12] over $W^{1,2}(\mathbb{R}^{2})$
@@
$\left(\int_{\mathbb{R}^{2}}(2P_{i})^{2}\mathrm{d}x\right)^{2}=\left(\int_{\mathbb{R}^{2}}\frac{|2P_{i}|}{1+|2P_{i}|}\left(1+|2P_{i}|\right)|2P_{i}|\mathrm{d}x\right)^{2}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\leq\int_{\mathbb{R}^{2}}\frac{\left(2P_{i}\right)^{2}}{\left(1+|2P_{i}|\right)^{2}}\mathrm{d}x\int_{\mathbb{R}^{2}}\left(1+|2P_{i}|\right)^{2}|2P_{i}|^{2}\mathrm{d}x$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\leq 4\int_{\mathbb{R}^{2}}\frac{\left(2P_{i}\right)^{2}}{\left(1+|2P_{i}|\right)^{2}}\mathrm{d}x\int_{\mathbb{R}^{2}}|2P_{i}|^{2}\mathrm{d}x\left(1+\int_{\mathbb{R}^{2}}|2\nabla P_{i}|^{2}\mathrm{d}x\right)$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\leq\frac{1}{2}\left(\int_{\mathbb{R}^{2}}|2P_{i}|^{2}\mathrm{d}x\right)^{2}+32\Bigg{\{}1+\left[\int_{\mathbb{R}^{2}}\frac{\left(2P_{i}\right)^{2}}{\left(1+|2P_{i}|\right)^{2}}\mathrm{d}x\right]^{4}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\left(\int_{\mathbb{R}^{2}}|2\nabla P_{i}|^{2}\mathrm{d}x\right)^{4}\Bigg{\}},~{}~{}i=1,2$ (4.32)
@@
$\left\|2P_{i}\right\|_{L^{2}(\mathbb{R}^{2})}\leq C\left[1+\int_{\mathbb{R}^{2}}\frac{\left(2P_{i}\right)^{2}}{\left(1+|2P_{i}|\right)^{2}}\mathrm{d}x+\int_{\mathbb{R}^{2}}|2\nabla P_{i}|^{2}\mathrm{d}x\right],~{}~{}i=1,2,$ (4.33)
@@
$F\left(w_{1},w_{2}\right)=\mathrm{e}^{2u_{2}^{0}}\left(\mathrm{e}^{2\left[\left(1-\frac{1}{2\alpha}\right)w_{1}+w_{2}\right]}-1\right)+\frac{2\alpha-1}{2\beta}\mathrm{e}^{2u_{1}^{0}}\left(\mathrm{e}^{2w_{1}}-1\right)+\frac{2\alpha-1}{\alpha\beta}\psi_{1}w_{1}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-\frac{\left(2\alpha-1\right)\left(\alpha+\beta\right)}{\alpha\beta}w_{1}+\frac{4\alpha}{\alpha+\beta}\psi_{2}w_{2}-2w_{2},$ (4.41)
@@ 4.2 Asymptotic behavior and quantized integrals
Finally, we estimate the decay rates for $u_{1},u_{2}$ and $|\nabla(mu_{1}+2u_{2})|,|\nabla(pu_{1}+qu_{2})|$, where $m,p,q$ are as defined by (3.18). For the coefficient matrix $A$ of the system (3.13), there exists a diagonal matrix $B$ such that the matrix $BA=M$ is positive definite and symmetric, where
@@
Therefore, we reduce (3.9)–(3.10) in $\mathbb{R}^{2}\backslash(0,0)$ to
@@
$\triangle g^{2}\geq\lambda_{0}g^{2}-C(x)g^{2},~{}~{}x\in\mathbb{R}^{2}\backslash(0,0),$ (4.50)
@@
$\triangle g^{2}\geq\left(1-\frac{\varepsilon}{2}\right)\lambda_{0}g^{2},~{}~{}x\in\mathbb{R}^{2}\backslash(0,0).$ (4.51)
@@
$~{}~{}~{}~{}\,\triangleq 2D\mathbf{h}+2D\left(E(x)-I\right)\mathbf{h}.$ (4.54)
@@
such that $TDT^{-1}=\Lambda=\mathrm{diag}\{\lambda_{3},\lambda_{4}\}$, where $\alpha=\dfrac{3}{2}-\dfrac{1}{2N}$ and
@@
$\triangle\mathbf{H}=2\Lambda\mathbf{H}+2TD\left(E(x)-I\right)T^{-1}\mathbf{H}.$ (4.56)
@@
$~{}~{}~{}~{}~{}~{}=4\mathbf{H}^{\tau}\Lambda\mathbf{H}+4\mathbf{H}^{\tau}TD\left(E(x)-I\right)T^{-1}\mathbf{H}$
$~{}~{}~{}~{}~{}~{}\geq\lambda\mathbf{|H|}^{2}-b(x)\mathbf{|H|}^{2},~{}~{}x\in\mathbb{R}^{2}\backslash(0,0),$ (4.57)
@@
We next are in a position to calculate the quantized integrals applying the decay estimates. Indeed, we have obtained the unique solution $(u_{1},u_{2})$ of the system (3.13) subject to the boundary condition (3.11) which vanish at infinity exponentially fast. Additionally, we deduce that $|\nabla(mu_{1}+2u_{2})|,|\nabla(pu_{1}+qu_{2})|$ all vanish at infinity at least as fast as $|x|^{-3}$, where $m,p,q$ are as defined by (3.18). Furthermore, from (4.1)–(4.2) and the exponential decay property of $|\nabla(mu_{1}+2u_{2})|$ and $|\nabla(pu_{1}+qu_{2})|$, we can see that $|\nabla(mP_{1}+2P_{2})|=O(|x|^{-3})$ and $|\nabla(pP_{1}+qP_{2})|=O(|x|^{-3})$ at infinity. Thus, according to the divergence theorem, we see that
$\int_{\mathbb{R}^{2}}\triangle\left(mP_{1}(x)+2P_{2}(x)\right)\mathrm{d}x=\int_{\mathbb{R}^{2}}\triangle\left(pP_{1}(x)+qP_{2}(x)\right)\mathrm{d}x=0.$ (4.60)
Now according to (4.3), (4.4), and (4.60) and the definitions of $\varphi_{i}(i=1,2)$, we can get the quantized integrals (3.1) as stated in Theorem 3.1.